# Fluctuation-Response Design Rules for Nonequilibrium Flows

### v1

@@ Abstract
This shows that imprecision couples to responsiveness: since the noise contribution scales with the squared sensitivity $(\partial\langle\psi\rangle/\partial k)^{2}$, a highly responsive step “amplifies” its own intrinsic noise into the overall variance. Crucially, this result moves beyond bounds like thermodynamic uncertainty relations [5, 9, 13] which constrain the magnitude of flux randomness based on global dynamical time irreversibility. Instead, Eq. (10) provides an exact mechanistic decomposition in terms of kinetic covariances and sensitivities. This approach directly identifies which specific molecular transitions act as the primary sources of the motor’s randomness, enabling the targeted dissection of the motor’s stochastic mechanism shown below.
@@
Universal symmetries and kinetic bounds. The geometric structure of the Jacobian A dictates universal principles that govern how any steady-state observables respond to perturbations in any ergodic network flow. From the identity $\textbf{A}^{-1}\textbf{A}=\textbf{I}$, we derive three sets of local response symmetries [1], valid for arbitrary CFT observables $\langle x\rangle\in\{\pi_{n},\tau_{ij},J_{c}\}$. This unifies and generalizes existing response relations [19, 3] within a single principle-based framework.
@@
where the right-hand side is a Kronecker delta, introducing an additional factor of 1 only if $\langle x\rangle=\pi_{n}$. This generalizes the response equalities derived by Owen et al. [19] from state probabilities to arbitrary CFT observables. On each edge $ij$, we have an Edge Reciprocity:
@@
Together with our covariance decomposition in Eq. (9), they underpin the flux fluctuation-response identities derived by Aslyamov et al. [3], while extending their validity to arbitrary CFT observables. On each cycle $c$, we find a Cycle Symmetry:
@@
This enforces a topological constraint: the cumulative sensitivity along any forward cycle must precisely balance that of its reversal (for observables $\langle x\rangle\neq J_{c}$). See [1] for derivation details and correspondences to past results.
Our framework also provides the physical origins of established response bounds. By identifying the algebraic coefficients $\Delta_{ij}$ and $\nabla_{ij}$ appeared in Aslyamov and Esposito [2] as specific rate sensitivities [1]:
@@
Code Availability: The code used to produce the Figures are available at https://doi.org/10.5281/zenodo.18611573.

### v2

# arXiv:2602.10957v2[cond-mat.stat-mech] 27 Feb 2026

@@ Abstract
This shows that imprecision couples to responsiveness: since the noise contribution scales with the squared sensitivity $(\partial\langle\psi\rangle/\partial k)^{2}$—a highly responsive step “amplifies” its own intrinsic noise into the overall variance. Crucially, this result moves beyond bounds like thermodynamic uncertainty relations [6, 10, 14] which constrain the magnitude of flux randomness based on global dynamical time irreversibility. Instead, Eq. (10) provides an exact mechanistic decomposition in terms of kinetic covariances and sensitivities. This approach directly identifies which specific molecular transitions act as the primary sources of the motor’s randomness, enabling the targeted dissection of the motor’s stochastic mechanism shown below.
@@
Universal symmetries and kinetic bounds. The geometric structure of the Jacobian A dictates universal principles that govern how any steady-state observables respond to perturbations in any ergodic network flow. From the identity $\textbf{A}^{-1}\textbf{A}=\textbf{I}$, we derive three sets of local response symmetries, valid for arbitrary CFT observables $\langle x\rangle\in\{\pi_{n},\tau_{ij},J_{c}\}$. This unifies and generalizes existing response relations [20, 3] within a single principle-based framework. See [1] for derivations.
@@
where the right-hand side is a Kronecker delta, introducing an additional factor of 1 only if $\langle x\rangle=\pi_{n}$. Physically, this corresponds to perturbing the site energy $E_{n}$ in an Arrhenius-like rate law ($k_{nj}\propto e^{E_{n}}$) [20], which tunes all outgoing rates simultaneously ($\partial_{E_{n}}\equiv\sum_{j(\neq n)}k_{nj}\partial_{k_{nj}}$). Eq. (11) thus generalizes the response equalities derived by Owen et al. [20] from state probabilities to arbitrary CFT observables. Applied to traffic or cycle fluxes, it reveals that these “node energy” perturbations scale all network fluxes proportional to local occupancy ($\partial_{E_{n}}\tau_{ij}=\pi_{n}\tau_{ij}$ and $\partial_{E_{n}}J_{c}=\pi_{n}J_{c}$).
On each edge $ij$, we have an Edge Reciprocity:
@@
Physically, this symmetry governs the barrier height $B_{ij}$ perturbation in an Arrhenius-like rate law ($k_{ij},k_{ji}\propto e^{-B_{ij}}$) [20], which tunes forward and backward rates symmetrically ($\partial_{B_{ij}}\equiv-k_{ij}\partial_{k_{ij}}-k_{ji}\partial_{k_{ji}}$). Eq. (12) reveals that the response for $\langle x\rangle\neq\tau_{ij}$ is strictly proportional to the net flux on the perturbed edge $J_{ij}\coloneqq\pi_{i}k_{ij}-\pi_{j}k_{ji}$: $\partial_{B_{ij}}\langle x\rangle=-(J_{ij}/\pi_{i})\partial_{k_{ij}}\langle x\rangle$. This implies that barrier perturbations cannot tune distributions $\pi_{n}$, cycle fluxes $J_{c}$, or other traffic terms $\tau_{uv}$ ($uv\neq ij)$ if the edge is “locally” at equilibrium ($J_{ij}=0$). They acquire control only through a nonzero nonequilibrium net flux $J_{ij}$. Together with our covariance decomposition in Eq. (9), these symmetries underpin the flux fluctuation-response identities derived by Aslyamov et al. [3], while extending their validity to arbitrary CFT observables.
On each cycle $c$, we find a Cycle Symmetry:
@@
This enforces a topological constraint: the cumulative sensitivity along any forward cycle must precisely balance that of its reversal (for observables $\langle x\rangle\neq J_{c}$, including all $\pi_{n},\tau_{ij},$ and the other cycle fluxes $J_{c^{\prime}}$).
Our framework also provides the physical origins of established response bounds. By identifying the algebraic coefficients $\Delta_{ij}$ and $\nabla_{ij}$ defined in Aslyamov and Esposito [2] as specific rate sensitivities [1]:
@@
Code Availability: The codes used to produce the Figures are available at https://doi.org/10.5281/zenodo.18611573.